Fermi National Laboratory


Accelerator Update for April 4, 2005 - April 6, 2005

Monday April 4
The day shift began with Operations monitoring the Recycler stash and the Antiproton Source (Pbar) stack, with the Tevatron (TeV) in shot setup, and with MiniBooNE taking beam. (Pbar stops stacking and Switchyard (SY) receives no beam during any shot setup.)

At 9:25 AM, Operations reported that a Main Injector (MI) crate had lost power for about an hour. It had backup power, but only for twenty minutes. Operations put shot setup on hold for twenty-five minutes while experts confirmed that the crate was okay.

At 10:39 AM, a Booster RF station (BRF16) tripped off for the first of two times in six minutes.

Operations established combination store 4075 at 11:02 AM with an initial luminosity of 111.74E30.

At 11:43 AM, BRF16 tripped off. It reset okay.

At 12:05 PM, MIRF15 tripped off for the first of three times in three minutes. It would not reset after the last trip. Operators contacted an expert.

Operators prepared MI-65 for access by NuMI experts at 2:51 PM. (NuMI was off due to a problem with its target.)

At 4:18 PM, Linac experts reported that the power amplifier tube for LRF5 was nearing the end of its useful service. If problems develop with this station, experts must be called.

RF experts reported at 5:11 PM that MIRF2 was operating normally.

At 5:20 PM, BRF16 tripped off. It reset okay.

Operations put the Recycler into shot setup to stash antiprotons at 5:48 PM.

At 5:50 PM, BRF16 tripped off for the first of two times in twenty-two minutes. It reset okay both times.

The Recycler transfer ended at 7 PM and Pbar resumed stacking.

Operations established beam to the Meson MCenter experiment at 7:13 PM.

At 9:13 PM, LRF5 tripped off. Operations called in experts.

At 9:39 PM, BRF8 tripped off. It reset okay.

Operations prepared Booster for access at 11:14 PM to check on sump #16. It was full and overflowing, but the water was going to other sump pits.

Tuesday April 5
The midnight shift began with Operations monitoring store 4075 and the stack and stash, and with beam off due to problems with LRF5.

Operations reported that LRF5 was back online at 12:33 AM and that Pbar had resumed stacking and MiniBooNE and MCenter had resumed taking beam.

At 1:56 AM, LRF3 tripped off. It reset okay.

Recycler began stashing antiprotons at 2 AM.

After the Recycler transfers ended, Pbar resumed stacking at 2:44 AM and MCenter resumed taking beam at 2:56 AM.

At 5:53 AM, BRF16 tripped off. It reset okay.

Linac experts turned off LRF5 at 9:26 AM to replace the power amplifier tube. This halted all downstream beam.

At 10:24 AM, a Booster quadrupole power supply (B:Q802) tripped off for the first of three times in twenty minutes. It reset okay each time.

At 10:51 Booster experts said that the 400 MeV area chopper (B:CHOP) had to have its "on" tube changed out. They began repairs.

Linac expert reported at 12:32 PM that they had completed their work on LRF5.

Booster experts reported at 1:34 PM that the chopper was fixed, but operators reported having trouble sending beam to Booster. Operations reported at 2:20 PM that the beam was going to the Linac dump rather than down the 400 MeV line. Experts continued to investigate.

Operations reported that beam through Linac and Booster had resumed by 2:53 PM. Pbar resumed stacking and MiniBooNE and MCenter resumed taking beam.

At 3:04 PM, LRF2 tripped off.

At 3:05 PM, Operations reported that the new chopper on tube was putting out noise and causing the "off" tube to fire. Experts investigated.

Pbar resumed stacking at 3:19 PM.

Operations reported at 4:09 PM that Booster experts added a new filter to the chopper.

At 5:15 PM, BRF16 tripped off. It reset okay.

At 5:31 PM, Operations reported that a Linac quadrupole power supply (L:QPS107) wouldn't regulate properly. Operators swapped control cards and had in back up by 5:54 PM.

At 7 PM, BRF3 tripped off. It reset okay.

Operations terminated store 4075 at 7:05 PM.

Operations put the TeV into shot setup at 7:25 PM.

At 8:32 PM, BRF16 tripped off. It reset okay.

At 9:02 PM, Operations reported that during the first transfer of antiprotons from Pbar, the TeV quenched at sector E4. The first transfer was supposed to hold 32mA, but it had held 41.4mA instead.

At 11:08 PM, BRF16 tripped off. It reset okay.

Operations put the TeV into a dry squeeze at 11:23 PM.

Operation put the TeV into shot setup at 11:52 PM.

Wednesday April 6
The midnight shift began with the TeV in shot setup, with Operations monitoring the stack and stash, and with MiniBooNE taking beam.

At 12:52 AM, a Booster power supply (DOG13) tripped due to a magnet water-flow switch. Operations contacted an expert who suggested that operators access the tunnel and convince the switch that it should resume working.

At 12:55 AM, Operations reported that an Accumulator septa (A:ISEP2V) had tripped off and wouldn't reset. Operations called in an expert who had the septa fixed by 1:52 AM.

Operators reported at 2:32 AM that they had accessed Booster, rapped on the flow switch for DOG13, and got it to reset.

Operations resumed shot setup at 2:44 AM.

At 3:02 AM, Operations reported that they'd been having trouble with the Accumulator kickers since the TeV quench. Operations contacted a Pbar expert for help who explained how to get Pbar back into its extraction mode.

Operations put the TeV through a dry squeeze at 3:42 AM.

Operations established combination store 4078 at 7:01 AM with an initial luminosity of 98.25E30. Pbar resumed stacking.

MCenter resumed taking beam at 7:40 AM.

The Plan for Wednesday
The plan for today is to monitor the store and to continue stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/6/2005   email Fermilab